OpenStudy (anonymous):

Solve the system using elimination. 5x + 4y = 12 3x – 3y = 18

OpenStudy (anonymous):

5x + 4y = 12 *3 3x – 3y = 18 *4

OpenStudy (anonymous):

so x=4 y=2

OpenStudy (anonymous):

36 72

OpenStudy (anonymous):

thanks